If the third term and fifth term of a geometric progression are 225 and 5625 respectively, determine the sixth term.
Answer by mananth(16946)

If the third term and fifth term of a geometric progression are 225 and 5625 respectively, determine the sixth term.
Tn = ar^(n01) is the n th term of a GP
T3= ar^(3-1)
T3=ar^2
225=ar^2

Similarly
t5=ar^4
5625=ar^4
divides ne equation by the other
225/5625= ar^2/ar^4
225/5625= 1/r^2
25/75= 1/r
r=3
plug r
225=ar^2
225=a*9
a=225/9
a=25
t6=ar^5
t6=25*3^5
t6=25*243
t6=6075
6th term = 6075